# Towards Identifiable Unsupervised Domain Translation: A Diversified Distribution Matching Approach

### v2

@@ 1 Introduction
$\bullet$ Simple Implementation via Auxiliary Variables. Our theoretical revelation naturally gives rise to a novel UDT learning criterion. This criterion aligns multiple pairs of conditional distributions across the source and target domains. We define these conditional distributions over (potentially overlapping) subareas of the source/target domains using auxiliary variables. We demonstrate that in practical applications such as unpaired I2I translation, obtaining these auxiliary variable-associated subsets can be a straightforward task, e.g., through available side information or querying the foundation models like CLIP (Radford et al., 2021). Consequently, our identifiability theory can be readily put into practice.
@@ 2 Preliminaries
In the context of domain translation, a linked $(\bm{x},\bm{y})$ pair can be regarded as cross-domain data samples that represent the same “content”, and the translation functions $(\bm{f}^{\star},\bm{g}^{\star})$ are responsible for changing their “appearances/styles”. The term “content” refers to the semantic information to be kept across domains after translation. In Fig. 1, the content is the digit (instead of the rotation); in Fig. 4 of Sec. 3, the content is related to the characteristics of the person. A closely related concept was used in representation learning; see (Von Kügelgen et al., 2021; Lyu et al., 2022).
@@
Supervised Domain Translation (SDT). In SDT, the corresponding pairs $(\bm{x},\bm{y})$ are assumed to be aligned a priori. Then, learning a translation function is essentially a regression problem—e.g., via finding $\bm{g}$ (or $\bm{f}$) such that $D(\bm{g}(\bm{x})||\bm{y})$ (or $D(\bm{f}(\bm{y})||\bm{x})$) is minimized over all given pairs, where $D(\cdot||\cdot)$ is any distance metric; see, e.g., (Isola et al., 2017; Wang et al., 2018).
@@
Lack of Translation Identifiability, MPA and Content Misalignment. Many works have noticed that distribution matching-type learning criterion may suffer from the lack of translation identifiability (Liu et al., 2017; Moriakov et al., 2020; Galanti et al., 2018b; 2021; Xu et al., 2022); i.e., the solution space of these criteria could have multiple solutions, and thus lack the ability to recover the ground-truth $\bm{g}^{\star}$ and $\bm{f}^{\star}$. The lack of identifiability often leads to issues such as content misalignment as we saw in Fig. 1.
To understand the identifiability challenge, let us formally define identifiability of any bi-directional UDT learning criterion:
Definition 1.
(Identifiability) Under the setting of Assumption 1, assume that $(\widehat{\bm{f}},\widehat{\bm{g}})$ is any optimal solution of a UDT learning criterion. Then, identifiability of $(\bm{f}^{\star},\bm{g}^{\star})$ holds under the UDT learning criterion if and only if $\widehat{\bm{f}}=\bm{f}^{\star}$ and $\widehat{\bm{g}}=\bm{g}^{\star}$ a.e.
@@ 3 Identifiable UDT via Diversified Distribution Matching
We call the proposed method diversified distribution matching for unsupervised domain translation (DIMENSION) 111 Note that we still use the term “unsupervised” despite the need of auxiliary information—as no paired samples are required. We avoided using “semi-supervised” or “weakly supervised” as these are often reserved for methods using some paired samples; see, e.g., (Wang et al., 2020; Mustafa & Mantiuk, 2020). . The following lemma shows that DIMENSION exactly realizes our idea in (6):
Lemma 1.
@@
Identfiiability Characterization. Lemma 1 means that solving the DIMENSION loss leads to conditional distribution matching as we hoped for in (6). Hower, it does not guarantee that $(\widehat{\bm{f}},\widehat{\bm{g}})$ found by DIMENSION satisfies $\widehat{\bm{f}}=\bm{f}^{\star}$ and $\widehat{\bm{g}}=\bm{g}^{\star}$. Towards establishing identifiability of the ground-truth translation functions via DIMENSION, we will use the following definition:
Definition 3 (Admissible MPA).
@@
Definition 4 (Sufficiently Diverse Condition (SDC)).
For any two disjoint sets $\mathcal{A},\mathcal{B}\subset\mathcal{X}$, where ${\cal A}$ and ${\cal B}$ are connected, open, and non-empty, there exists a ${u}_{(\mathcal{A},\mathcal{B})}\in\{u_{1},\dots,u_{I}\}$ such that $\mathbb{P}_{\bm{x}|u={u}_{(\mathcal{A},\mathcal{B})}}[\mathcal{A}]\neq\mathbb{P}_{\bm{x}|u={u}_{(\mathcal{A},\mathcal{B})}}[\mathcal{B}].$ Then, the set of conditional distributions $\{\mathbb{P}_{\bm{x}|u=u_{i}}\}_{i=1}^{I}$ is called sufficiently diverse.
Definition 4 puts the desired “diversity” into context. It is important to note that the SDC only requires the existence of a certain ${u}_{(\mathcal{A},\mathcal{B})}\in\{u_{1},\ldots,u_{I}\}$ for a given disjoint set pair $({\cal A},{\cal B})$. It does not require a unified ${u}$ for all pairs; i.e., $u_{(\mathcal{A},\mathcal{B})}$ needs not to be the same as $u_{({\mathcal{A}^{\prime},\mathcal{B}^{\prime}})}$ for $(\mathcal{A},\mathcal{B})\neq(\mathcal{A}^{\prime},\mathcal{B}^{\prime})$. Fig. 6 shows a simple example where the two conditional distributions satisfy the SDC. In more general cases, this implies that if the PDFs of the conditional distributions exhibit different “shapes” over their supports, SDC is likely to hold. Using SDC, we show the following translation identifiability result:
Theorem 1 (Identifiability).
Suppose that Assumption 1 holds. Let ${\sf E}_{i,j}$ denote the event that the pair $(\mathbb{P}_{\bm{x}|u=u_{i}},\mathbb{P}_{\bm{x}|u=u_{j}})$ does not satisfy the SDC. Assume that ${\rm Pr}[{\sf E}_{i,j}]\leq\rho$ for any $i\neq j$, where $i,j\in[I]$. Let $(\widehat{\bm{f}},\widehat{\bm{g}})$ be from an optimal solution of the DIMENSION loss (7). Then, there is no admissible MPA of $\{\mathbb{P}_{\bm{x}|u=u_{i}}\}_{i=1}^{I}$ of the solution, i.e., $\widehat{\bm{f}}=\bm{f}^{\star},~{}a.e.$ and $\widehat{\bm{g}}=\bm{g}^{\star},~{}a.e.$ with a probability of at least $1-\rho^{I\choose 2}$.
@@
Definition 5 (Relaxed Condition: $r$-SDC).
@@
Theorem 2 (Robust Identifiability).
@@ 5 Numerical Validation
MNIST to Rotated MNIST. Fig. 7 shows the results. In this case, we use $u\in\{1,\ldots,10\}$, i.e., the labels of the identity of digits, as the alphabet of the auxiliary variable. Note that knowing such labels does not mean that the cross-domain pairs $(\bm{x},\bm{y})$ are known. One can see that DIMENSION learns translates the digits to their corresponding rotated versions. But the baselines sometimes misalign the samples. The results are consistent with our analysis (see Sec. F.6 for more results).
@@ A.1 Notation
  • For a (random) vector $\bm{x}$, $x(i)$ and $[\bm{x}]_{i}$ denote the $i$th element of $\bm{x}$, and $\bm{x}(i:j)$ denotes $[x(i),x(i+1),\dots,x(j)]$.
@@ Appendix B Proof of Lemmas and Facts
$[\bm{d}(\overline{\bm{x}})]_{i}:=F\left(\overline{x}(i)~{}\big{|}~{}\bm{x}(1:i-1)=\overline{\bm{x}}(1:i-1)\right),\quad i=1,\dots,N,$
@@
${\color[rgb]{0.1953125,0.66796875,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.66796875,0.1953125}\pgfsys@color@rgb@stroke{0.1953125}{0.66796875}{0.1953125}\pgfsys@color@rgb@fill{0.1953125}{0.66796875}{0.1953125}\bm{h}_{U}(\bm{z})=-\bm{z}+2\mu},$
where $\bm{\mu}=[\nicefrac{{1}}{{2}},\dots,\nicefrac{{1}}{{2}}]^{\!\top\!}\in\mathbb{R}^{D_{x}}$. This concludes the proof. ∎
Lemma 1.
@@
Hence, there exists an admissible MPA of $\{\mathbb{P}_{\bm{x}|u=u_{i}}\}_{i=1}^{I}$ if and only if there exists an admissible MPA of $\{\mathbb{P}_{\bm{y}|u=u_{i}}\}_{i=1}^{I}$. ∎
@@ C.1 Proof of Theorem 1
Theorem 1.
Suppose that Assumption 1 holds. Let ${\sf E}_{i,j}$ denote the event that the set $\{\mathbb{P}_{\bm{x}|u=u_{i}},\mathbb{P}_{\bm{x}|u=u_{j}}\}$ does not satisfy the SDC. Assume that ${\rm Pr}[{\sf E}_{i,j}]\leq\rho$ for any $i\neq j$, where $i,j\in[I]$. Let $(\widehat{\bm{f}},\widehat{\bm{g}})$ be from an optimal solution of the DIMENSION loss (7). Then, there is no admissible MPA of $\{\mathbb{P}_{\bm{x}|u=u_{i}}\}_{i=1}^{I}$ of the solution, i.e., $\widehat{\bm{f}}=\bm{f}^{\star},~{}a.e.,\quad\widehat{\bm{g}}=\bm{g}^{\star},~{}a.e.,$ with a probability of at least $1-\rho^{I\choose 2}$.
@@
Therefore, we always have $\widehat{\bm{g}}=\bm{g}^{\star}$, a.e. By role symmetry of $\bm{f}^{\star}$ and $\bm{g}^{\star}$ (also see Fact 2), we also have $\widehat{\bm{f}}=\bm{f}^{\star}$, a.e. ∎
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\bigcap_{i,j\in[I],i<j}{\rm Pr}[{\sf E}_{i,j}]$
@@
Hence, $\{\mathbb{P}_{\bm{x}|u=u_{i}}\}_{i=1}^{I}$ are sufficiently diverse with probability at least $1-\rho^{I\choose 2}$, which implies that $\widehat{\bm{f}}=\bm{f}^{\star}$ and $\widehat{\bm{g}}=\bm{g}^{\star}$ with probability at least $1-\rho^{I\choose 2}$. ∎
@@ C.2 Proof of Theorem 2
Theorem 2 (Robust Identifiability).
@@
Proof Idea: The main idea behind the proof of Lemma A.2 is to first note that any point outside of ${\rm cl}(\mathcal{E}_{\bm{h}})$ is stationary under the transformation $\bm{h}$ (i.e., $\bm{h}(\bm{x})=\bm{x}$). Next, if there was a point $\overline{\bm{x}}$ from a connected component $\mathcal{C}_{1}\in{\rm conn}({\rm cl}(\mathcal{E}_{h}))$ was such that $\bm{h}(\overline{\bm{x}})$ was not in $\mathcal{C}_{1}$, then it should be either in $\mathcal{X}\backslash{\rm cl}(\mathcal{E}_{\bm{h}})$ or in ${\rm cl}(\mathcal{E}_{\bm{h}})\backslash\mathcal{C}_{1}$. However, since $\bm{h}$ is invertible, $\bm{h}$ cannot map a point from $\mathcal{C}_{1}$ to a $\mathcal{X}\backslash{\rm cl}(\mathcal{E}_{\bm{h}})$. Therefore $\bm{h}(\overline{\bm{x}})$ should lie in $\mathcal{E}_{h}\backslash\mathcal{C}_{1}$. But this will make the function $\bm{h}$ discontinuous. Hence, $\bm{h}(\overline{\bm{x}})$ should be in $\mathcal{C}_{1}$.
@@
$\bm{h}(\widetilde{\bm{x}})=\widetilde{\bm{x}},~{}\forall\widetilde{\bm{x}}\in\mathcal{X}\backslash{\rm cl}(\mathcal{E}_{h})$ $\stackrel{{\scriptstyle(a)}}{{\implies}}\bm{h}(\overline{\bm{x}})\neq\widetilde{\bm{x}},~{}\forall\widetilde{\bm{x}}\in\mathcal{X}\backslash{\rm cl}(\mathcal{E}_{h})$
@@
Because of the continuity of $\bm{h}$, the set $\bm{h}(\mathcal{C}_{1})$ is a closed connected set containing $\bm{h}(\overline{\bm{x}})$. However, $\bm{h}(\mathcal{C}_{1})\cap(\mathcal{X}\backslash{\rm cl}(\mathcal{E}_{\bm{h}}))=\phi$. This means that
@@
However, take any point $\bm{z}\in\mathcal{N}_{\delta}(\bm{x}_{b})\backslash{\rm cl}(\mathcal{E}_{\bm{h}})$ with $\|\bm{z}-\bm{x}_{b}\|_{2}<\epsilon/4$. Such a point exists because any neighborhood of a point on the boundary of a closed set has a non-empty intersection with the complement of the closed set. Therefore, we have
@@
[…]
This concludes the proof. ∎
Lemma A.3.
@@
Note that such a $\bm{w}$ has to exist. Suppose that such a $\bm{w}$ does not exist. Then, ${\rm cl}(\mathcal{R}_{d})\cap{\rm bd}(\mathcal{N}_{d})=\phi$, which means that $\mathcal{R}_{d}$ would be disconnected from $\overline{\mathcal{C}}\backslash\mathcal{N}_{d}$. By the definition of $\mathcal{R}_{d}$, $\mathcal{R}_{d}$ is then disconnected from $\mathcal{T}_{d}\backslash\mathcal{R}_{d}$, which implies that $\overline{\mathcal{C}}$—that is a union of $\mathcal{R}_{d}$, $\mathcal{T}_{d}\backslash\mathcal{R}_{d}$, and $\overline{\mathcal{C}}\backslash\mathcal{N}_{d}$—is disconnected. This is a contradiction. Hence ${\rm cl}(\mathcal{R}_{d})\cap{\rm bd}(\mathcal{N}_{d})\neq\phi$ holds.
@@
$30<{\rm dist}(\mathcal{R}_{d},\overline{\bm{h}}(\mathcal{R}_{d}))<\epsilon/$ 3 .
@@
This concludes the proof. ∎
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\bigcap_{i,j\in[I],i<j}{\rm Pr}[{M}_{i,j}]$
@@
The same result follows for $\widehat{\bm{f}}$ if $\bm{f}^{\star}$ is $L$-Lipschitz continuous, following the same procedure as above. ∎
@@ Appendix E Synthetic Data Experiments
Figure 14: TE under various $I$’s.
@@ F.2 Neural Network Details
UpSample Upsample using nearest neighbor with scale factor of $2$
@@
K-$N$ Kernel (filter) of size $N$
S-$N$ Stride of size $N$
ZP-$N$ Zero Padding of size $N$
C-$M$-$N$ $M$ input and $N$ output channels
The translation neural networks, $\bm{g}$ and $\bm{f}$, for images of size $256\times 256$ follow the architecture outlined in Table 3. For images of size $128\times 128$, a modified architecture is used, where one down-sampling layer (see Layer ##\##6) and one up-sampling layer (Layer #$\#11$) in Table 3 are not included. For images of size $32\times 32$, three down-sampling layers (indices from #$\#4$ to #$\#6$) and three up-sampling layers (indices from #$\#11$ to #$\#13$) are not included.

### v3

@@ 1 Introduction
$\bullet$ Simple Implementation via Auxiliary Variables. Our theoretical revelation naturally gives rise to a novel UDT learning criterion. This criterion aligns multiple pairs of conditional distributions across the source and target domains. We define these conditional distributions over (overlapping) sub-domaions of the source/target domains using auxiliary variables. We demonstrate that in practical applications such as unpaired I2I translation, obtaining these sub-domains can be a straightforward task, e.g., through available side information or querying the foundation models like CLIP (Radford et al., 2021). Consequently, our identification theory can be readily put into practice.
@@ 2 Preliminaries
In the context of domain translation, a linked $(\bm{x},\bm{y})$ pair can be regarded as cross-domain data samples that represent the same “content”, and the translation functions $(\bm{f}^{\star},\bm{g}^{\star})$ are responsible for changing their “appearances/styles”. The term “content” refers to the semantic information to be kept across domains after translation. In Fig. 1, the content is the identity of the digit ( other than writing style or the rotation); in Fig. 4 of Sec. 3, the content can be understood as the shared characteristics of the person in both the cartoon and the photo domains, which can collectively identify the person.
@@
Supervised Domain Translation (SDT). In SDT, the corresponding pairs $(\bm{x},\bm{y})$ are assumed to be aligned a priori. Then, learning a translation function is essentially a regression problem—e.g., via finding $\bm{g}$ (or $\bm{f}$) such that $D(\bm{g}(\bm{x})||\bm{y})$ (or $D(\bm{f}(\bm{y})||\bm{x})$) is minimized over all given pairs, where $D(\cdot||\cdot)$ is a certain “distance” measure; see, e.g., (Isola et al., 2017; Wang et al., 2018).
@@
Lack of Translation Identifiability, MPA and Content Misalignment. Many works have noticed that distribution matching-type learning criterion may suffer from the lack of translation identifiability (Liu et al., 2017; Moriakov et al., 2020; Galanti et al., 2018b; 2021; Xu et al., 2022); i.e., the solution space of these criteria could have multiple solutions, and thus lack the ability to recover the ground-truth $\bm{g}^{\star}$ and $\bm{f}^{\star}$. The lack of identifiability often leads to issues such as content misalignment as we saw in Fig. 1. To understand the identifiability challenge, let us formally define identifiability of any bi-directional UDT learning criterion:
Definition 1
[…]
@@ 3 Identifiable UDT via Diversified Distribution Matching
We call the proposed method diversified distribution matching for unsupervised domain translation (DIMENSION)  [^1] Note that we still use the term “unsupervised” despite the need of auxiliary information—as no paired samples are required. We avoided using “semi-supervised” or “weakly supervised” as these are often reserved for methods using some paired samples; see, e.g., (Wang et al., 2020; Mustafa & Mantiuk, 2020). . The following lemma shows that DIMENSION exactly realizes our idea in (6):
Lemma 1
@@
[…]
Definition 3 (Admissible MPA)
@@
Definition 4 (Sufficiently Diverse Condition (SDC))
[…]
Definition 4 puts the desired “diversity” into context. It is important to note that the SDC only requires the existence of a certain ${u}_{(\mathcal{A},\mathcal{B})}\in\{u_{1},\ldots,u_{I}\}$ for a given disjoint set pair $({\cal A},{\cal B})$. It does not require a unified u{u} for all pairs; i.e., $u_{(\mathcal{A},\mathcal{B})}$ needs not to be the same as $u_{({\mathcal{A}^{\prime},\mathcal{B}^{\prime}})}$ for $(\mathcal{A},\mathcal{B})\neq(\mathcal{A}^{\prime},\mathcal{B}^{\prime})$. Fig. 6 shows a simple example where the two conditional distributions satisfy the SDC. In more general cases, this implies that if the PDFs of the conditional distributions exhibit different “shapes” over their supports, SDC is likely to hold. Using SDC, we show the following translation identifiability result:
Theorem 1 (Identifiability)
[…]
@@
Definition 5 (Relaxed Condition: $r$-SDC)
@@
Theorem 2 (Robust Identifiability)
@@ 5 Numerical Validation
MNIST to Rotated MNIST. Fig. 7 shows the results. In this case, we use $u\in\{1,\ldots,10\}$, i.e., the labels of the identity of digits, as the alphabet of the auxiliary variable. Note that knowing such labels does not mean that the cross-domain pairs $(\bm{x},\bm{y})$ are known. Alternatively, one can also use digit shapes as the alphabets (see Sec. F.6). One can see that DIMENSION learns to translate the digits to their corresponding rotated versions. But the baselines sometimes misalign the samples. The results are consistent with our analysis (see Sec. F.6 for more results).
@@ 6 Conclusion
Acknowledgement. This work is supported in part by the Army Research Office (ARO) under Project ARO W911NF-21-1-0227, and in part by the National Science Foundation (NSF) CAREER Award ECCS-2144889.
@@ A.1 Notation
  • For a (random) vector $\bm{x}$, $x(i)$ and $\bm{x}]_{i}$ denote the $i$th element of $\bm{x}$, and $\bm{x}(i:j)$ denotes $x(i),x(i+1),\dots,x(j)]$.
@@ Appendix B Proof of Lemmas and Facts
$\bm{d}(\overline{\bm{x}})]_{i}:=F\left(\overline{x}(i)~\big{|}~\bm{x}(1:i-1)=\overline{\bm{x}}(1:i-1)\right),\quad i=1,\dots,N,$
@@
$\color[rgb]{0.1953125,0.66796875,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.66796875,0.1953125}\bm{h}_{U}(\bm{z})=-\bm{z}+2\mu$},
where $\bm{\mu}=[\nicefrac{{1}}{{2}},\dots,\nicefrac{{1}}{{2}}]^{\!\top\!}\in\mathbb{R}^{D_{x}}$. This concludes the proof.
nobreak  ∎
Lemma 1
@@
Hence, there exists an admissible MPA of $\{\mathbb{P}_{\bm{x}|u=u_{i}}\}_{i=1}^{I}$ if and only if there exists an admissible MPA of $\{\mathbb{P}_{\bm{y}|u=u_{i}}\}_{i=1}^{I}$.
nobreak  ∎
@@ C.1 Proof of Theorem 1
Theorem 1
[…]
@@
Therefore, we always have $\widehat{\bm{g}}=\bm{g}^{\star}$, a.e. By role symmetry of $\bm{f}^{\star}$ and $\bm{g}^{\star}$ (also see Fact 2), we also have $\widehat{\bm{f}}=\bm{f}^{\star}$, a.e.
nobreak  ∎
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\bigcap_{i,j\in[I],i<j}{\rm Pr}[{\sf E}_{i,j}]$
@@
Hence, $\{\mathbb{P}_{\bm{x}|u=u_{i}}\}_{i=1}^{I}$ are sufficiently diverse with probability at least $1-\rho^{I\choose 2}$, which implies that $\widehat{\bm{f}}=\bm{f}^{\star}$ and $\widehat{\bm{g}}=\bm{g}^{\star}$ with probability at least $1-\rho^{I\choose 2}$.
nobreak  ∎
@@ C.2 Proof of Theorem 2
Theorem 2 (Robust Identifiability)
@@
[…]
@@
[…]
@@
[…]
@@
$\bm{x}\in\mathcal{C}\text{ for any }\mathcal{C}\in{\rm conn}({\rm cl}(\mathcal{E}_{\bm{h}}))\implies\bm{h}(\bm{x})\in\mathcal{C}.$
This concludes the proof.
nobreak  ∎
Lemma A.3
@@
[…]
@@
$0<{\rm dist}(\mathcal{R}_{d},\overline{\bm{h}}(\mathcal{R}_{d}))<\epsilon/3.$
@@
This concludes the proof.
nobreak  ∎
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\bigcap_{i,j\in[I],i<j}{\rm Pr}[{M}_{i,j}]$
@@
The same result follows for $\widehat{\bm{f}}$ if $\bm{f}^{\star}$ is $L$-Lipschitz continuous, following the same procedure as above.
nobreak  ∎
@@ Appendix E Synthetic Data Experiments
Figure 14: TE under various II’s.
@@ F.2 Neural Network Details
UpSample Upsample using nearest neighbor with scale factor of 22
@@
K-NN Kernel (filter) of size NN
S-NN Stride of size NN
ZP-NN Zero Padding of size NN
C-$M$-NN $M$ input and $N$ output channels
The translation neural networks, $\bm{g}$ and $\bm{f}$, for images of size $256\times 256$ follow the architecture outlined in Table 3. For images of size $128\times 128$, a modified architecture is used, where one down-sampling layer (see Layer #\#6) and one up-sampling layer (Layer #11\#11) in Table 3 are not included. For images of size $32\times 32$, three down-sampling layers (indices from #4\#4 to #6\#6) and three up-sampling layers (indices from #11\#11 to #13\#13) are not included.